OpenStudy (anonymous):

Rationalize the Denominator: 3x/√5:(x^(2)yz^(4))

OpenStudy (anonymous):

what does the colon represent?

OpenStudy (anonymous):

Ugh i formated it on the front page, it should be the fith root of x^2yz^4

OpenStudy (anonymous):

\[\sqrt[5]{x^2yz^4}\]

OpenStudy (anonymous):

ah ok, well the first rule of these equations is to get rid of the radical in the denominator

OpenStudy (anonymous):

Ok, by mutliplying into the denom?

OpenStudy (anonymous):

since it is the fifth route, you need to multiply until each variable has a factor of 5 as an exponent

OpenStudy (anonymous):

So, x10y5z20?

OpenStudy (anonymous):

ah ok, well the first rule of these equations is to get rid of the radical in the denominator

OpenStudy (anonymous):

What about the x by the three?

OpenStudy (anonymous):

Oh ok. Sorry im a bit laggy

OpenStudy (anonymous):

\[\sqrt[5]{x ^{3}y ^{4}z}\]

OpenStudy (anonymous):

if you multiply that to the denominator you will get \[\sqrt[5]{x ^{5}y ^{5}z ^{5}}\] or xyz

OpenStudy (anonymous):

Im confused?

OpenStudy (anonymous):

\[3x/\sqrt[5]{x ^{2}yz ^{3}} \times \sqrt[5]{x ^{3}y ^{4}z} / \sqrt[5]{x ^{3}y ^{4}z}\]

OpenStudy (anonymous):

then it becomes \[(3x \sqrt[5]{x ^{3}y ^{4}z}) / xyz \]

OpenStudy (anonymous):

The z is to the four not the y

OpenStudy (anonymous):

ok, im sorry for the confusion, this is my first time trying this

OpenStudy (anonymous):

is this problem, you have a radical in the denominator, to get rid of that you must make each exponent underneath the radical 5

OpenStudy (anonymous):

that way the exponent and the radical cancel out